OpenStudy (maaz):

If f(x) = sin(x) for all x, then the average value of f on the interval [0, π] is (A) ½ (B) 1/π (C) π/2 (D) 2/pie

ganeshie8 (ganeshie8):

\[1/(\pi -0)\int\limits_{0}^{\pi} sinx\]